Can you do this for me please
5[2-3(6x-1)]
Im not sure how you do this

Found 2 solutions by Earlsdon, Theo:
Answer by Earlsdon(6294)

Simplify(?):
5[2-3(6x-1)] First, apply the distributive property to the parentheses inside the brackets.
5[2-3*6x-(-3)*(1)] Simplify the inside of the brackets.
5[2-18x+3] = 5[5-18x] Now apply the distributive property again.
5*5-5*18x = 25-90x

Answer by Theo(13342)

5[2-3(6x-1)]
you need to resolve -3(6x-1) first.
that becomes -3 * 6x - 3*-1 which becomes:
-18x + 3
your equation becomes:
5(2-18x+3)
combine like terms to get:
5(-18x+5)
simplify to get:
-90x + 25
that should be your answer.

to confirm that this answer is good, let x be any value that's easy to calculate and then solve the original equation and the final equation using that value of x.

if the answer from the original equation is the same as the answer from the final equation, then you did good.

in this case, i'll let x = 2.
the original equation is:
5(2-3(6x-1))
replace x with 2 and the equation becomes:
5(2-3(6*2-1)) which becomes:
5(2-3(12-1)) which becomes:
5(2-3(11) which becomes:
5(2-33) which becomes:
5(-31) which becomes:
-155

the final equation is:
-90x+25
replace x with 2 to get:
-90*2+25
simplify to get:
-180+25
simplify further to get:
-155

answers from the original equation and the final equation are the same so you can reasonably assume that the answer is correct.
final equation is -90x+24